WITH WHAT VELOCITY SHOULD AN ALPHA PARTICLE TRAVEL TOWARDS THE NUCLEUS OF A COPPER ATOM SO AS TO ARRIVE AT A DISTANCE 1/10000000000000 METRE FROM THE NUCLEUS OF THE COPPER ATOM


 

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

let m be mass of alpha particle and M be mass of copper nucleus. Let 'v' be the required velocity.

mv2/2=(1/4πε0)*mM/r, where r=1/10000000000000 METRE